# Optomechanical Systems with Linear and Quadratic Position Couplings: Dynamics and Optimal Estimation

### v1

@@ IV Results
We start our analysis by considering a simple scenario in which the initial photonic state $\hat{\sigma}(0)$ can be written on the Hilbert space spanned by the zero-photon state $|0\rangle$ and the one-photon state $|1\rangle$. Thus, $\hat{\sigma}(0)$ is a $2\times 2$ positive semi-definite density matrix that can be written as
$\hat{\sigma}(0)=\frac{1}{2}(\mathbf{I}+r_{x}\sigma_{x}+r_{y}\sigma_{y}+r_{z}\sigma_{z}),$ (44)
@@
where $\sigma_{i}$ with $i=x,y,z$ are the Pauli matrices and the coefficients $r_{i}$ are real numbers s.t. $r_{x}^{2}+r_{y}^{2}+r_{z}^{2}\leq 1$, with the equality satisfied when $\hat{\sigma}(0)$ is a pure state.
IV.1 Unitless Model
To present the system dynamics in an intuitive manner, we first consider a unitless model in which all relevant physical parameters are set to one and unitless, while the linear coupling coefficient $g_{1}$ is set to zero and the quadratic $g_{2}$ to 0.01. See Table 1.
UM RS
[…]
$n_{th}$ $1$ $10^{5}$
[…]
$g_{2}$ $0.01$ $2\pi\times 4.46\times 10^{24}$ $[\text{Hz}/\text{m}^{2}]$
Table 1: Table of parameter values used in the unitless model (UM) and the real-world system (RS). SM values are unitless; RS values are in SI units.
In this section, we consider either the pure state
@@ IV.1 Unitless Model
both parametrized by the single parameter $s$. Note that the coefficients $r_{x}$ and $r_{y}$ of the pure states $|\psi_{s}\rangle$ and $|\psi_{1-s}\rangle$ coincide. In the following, we assign the value $s=0.2$, while a similar dynamics has been observed for different choices of $s.$
IV.1.1 System dynamics
Figure 1: The trajectory in the X-Y plane of the Bloch sphere of the evolution of the quantum states (45),(46) parameterized by $s=0.1$ (red) and $s=0.2$(blue). The system is evolved from time is 0 to $4\pi/\delta,$ which contains two complete cycles of the fast dynamic oscillation. The first cycles are denoted by solid lines and the second by dotted lines. To make the spiraling evolution more visible, the values of Bloch sphere coordinates $r_{x}$ and $r_{y}$ are scaled by $r^{2}=r_{x}^{2}+r_{y}^{2}.$ The evolution depends only on $s$, regardless on the purity of the state.
In Fig. 1 we show the parametric evolution of the optical state (33) expressed in terms of the coefficients $r_{x}$ and $r_{y}$, allowing the system evolve between $t=0$ and $t=\frac{4\pi}{\delta},$ with $\delta=\omega_{1}-\omega_{0}.$ The evolution is shown for two choices of the parameter $s$ of Eqs. (45) and (46). In fact, by solving the integral in Eq. (33), we can see that the diagonal elements of the optical density matrix are time-independent and therefore $r_{z}$ is a constant. Thus, the projection on the X-Y plane of the Bloch sphere suffices to describe the evolution of the optical system, while its value is independent from the purity $\mathcal{P}$ of the system, and can be parametrized solely by $s$. Note that, in order to make the features of the trajectories more visible, the axes have been scaled accordingly to $r_{x}/r^{2}$ and $r_{y}/r^{2}$, where $r^{2}=r_{x}^{2}+r_{y}^{2}.$ Comparing the first cycle in X-Y plane to the second, one observes the inward spiral of the coordinates towards the center of the Bloch sphere, which represents the maximally mixed state. Thus, the inward spiral shows a decrease in purity. Furthermore, we show in Figure 2 the evolution of the $r_{x}$ component. After about 50 fast cycles, the system spirals into the innermost point corresponding to the lowest purity, and then spirals outward again until it is back at the initial, outermost coordinates corresponding to the highest purity. We plot the time evolution of the purity $\mathcal{P}$ in Fig. 3 for a pure and mixed state with $s=0.2$.
@@ IV.1.1 System dynamics
Figure 2: Time evolution of $r_{x}$, the $x$ coordinate of the system in Bloch sphere, here plotted for the initial pure state $|\psi_{s}\rangle$ with $s=0.2$. Time is labeled by periods of complete fast oscillation, two of which are shown in Figure 1. Enveloping the fast oscillation is the beating oscillation $\Delta=\Omega_{1}-\Omega$, roughly $100$ times slower than the fast frequency.
Figure 3: The time evolution of the purity $\mathcal{P}$ for the pure (blue continuous line) and the mixed (red dashed line) states defined in Eqs. (45),(46) respectively, with $s=0.2$. Time is scaled by $\frac{2\pi}{\Delta}$, the natural oscillation period. In this figure, four oscillation periods have been plotted.
IV.1.2 Spectral Analysis
The time dependence of the optical coefficients as expressed in Eq. (33) is given by the term $\det(A)$, with main frequency being the beat frequency $\Delta=\Omega_{1}-\Omega$, and by the $\exp(i[\Phi_{1}(t)-\Phi_{0}(t)])$ factor, which contains the faster frequency $\delta=\omega_{1}-\omega_{0}$, while the last term $\exp(\frac{1}{2}b^{T}A^{-1}b+c)$ is a time-independent constant. In order to reveal the underlying frequencies, we performed a Fourier Transform on the dynamics of the $r_{x}$ coefficient,
[…]
which captures the off-diagonal time changing components of the optical density matrix. The two frequencies with larger amplitude are $\Delta$ and $\delta=\omega_{c}+\Delta/2$, while successive frequencies have an exponentially decaying amplitude, as shown in Figs. 4 and  5 respectively.
Figure 4: The absolute value of the Fourier transform of the coefficient $r_{x}(t)$, plotted in the range of the slow frequencies. The peaks are approximately multiples of $\frac{1}{3}\frac{\Delta}{2\pi}$.
Figure 5: The absolute value of the Fourier transform of the coefficient $r_{x}(t)$, plotted in the range of the fast frequencies. The first peaks is $\frac{\delta}{2\pi}$. This term originates from the $\exp(i[\Phi_{1}-\Phi_{0}]),$ whose exponent is proportional to $\omega_{1}-\omega_{0}$ in our system with $g_{1}=0.$ The peaks display an exponential decay in relative amplitude.
Looking at Fig. 4, we see that the major peaks are multiples of the $32\pi\Delta/$ 3. Conversely, the peaks in Fig. 5 represent the different periods of revolution of the $r_{x}$ coefficient, as can be also observed in Fig. 1, by comparing the position of the two $t=2\pi/\delta$ evolutions.
IV.1.3 Quantum Fisher Information
In Fig. 6 we show the QFI of the pure state $|\psi_{s}\rangle$ parameterized by $s=0.2$ numerically calculated using the definition (37). Note that a similar behavior is observed for other choices of the parameter $s$. The evolution of the QFI of the pure state presents sharp peaks at $t=2\pi/\Delta$. In Fig. 7, we show the temporal evolution of the QFI for the initial mixed state (46) with $s=0.2$. We calculated the QFI using the simplified analytic expression for $L_{g}$ given in Equation (40). Even for an initial mixed state, the periodicity of the peaks of the QFI coincide with the periodicity of the purity of the system. Around the peaks, the off-diagonal elements of the density matrix show greater sensitivity to changes of the $g_{2}$ parameter, as can be verified by solving Eq. (33). Contrary to the pure state, the peaks for the mixed state present a fast oscillation as we describe next.
Figure 6: Quantum Fisher Information as a function of time for the pure state with $s=0.2$, calculated with numerical integration. Periodic peaks are observed at the regular time interval of $2\pi/\Delta,$ entirely overlapping with the evolution of the purity.
@@ IV.1.3 Quantum Fisher Information
Figure 7: Quantum Fisher Information as a function of time for the mixed state with $s=0.2$, calculated analytically. Periodic peaks are observed at the regular time interval of $2\pi/\Delta.$
IV.1.4 Classical Fisher Information
@@
By solving Eq. (33), we see that the diagonal elements remain time-independent throughout the system evolution. Thus, the CFI calculated from the photodetection PVM, Eq. (41) is null, as information can be derived from the dynamics of the off-diagonal elements only. Conversely, the set of PVM related to the BHD, as defined in Eq. (43) provide information about the off-diagonal elements, and they are chosen to calculate the CFI. When the phase $\phi=0$ they reduce to
$\Pi_{0}=|+\rangle\langle+|,\;\Pi_{1}=|-\rangle\langle-|,$
@@
where |±$|\pm\rangle=\frac{1}{\sqrt{2}}(|0\rangle\pm|1\rangle)$, which lead to the measurement probability distribution:
$P_{0}=\frac{1}{2}(1+\rho_{01}+\rho_{10}),\;P_{1}=\frac{1}{2}(1-\rho_{01}-\rho_{10}).$
Note that $\frac{1}{2}(\rho_{01}+\rho_{10})$ is simply the real part of both the off-diagonal elements.
@@
We compare the CFI obtained with the above described set of PVM to the QFI in Fig. 8, where we show the result around the first QFI peak set at $t\approx\Delta/2\pi$ in the case of an initial mixed state with $s=0.2$. The CFI is fast oscillating, with period $2\pi/\delta$ and its peak is slightly shifted with respect to the QFI peak. However, we find that when the CFI is maximum, its value corresponds to the QFI value. Similar behavior is obtained both for different choices of $s$ and for an initial pure state.
@@
Figure 8: Classical Fisher Information (black dashed line) and Quantum Fisher Information (red continuous line) as a function of time zoomed in around the first period of the system’s beating oscillation, i.e., $t=\frac{2\pi}{\Delta}$.
IV.2 Real-world system
We take the numerical values from the paper [35], which describes a suspended SiN membrane that oscillates with frequency $\Omega=2\pi\times 134\text{ kHz}$ placed in the middle of a Fabry-Pérot optical cavity of length $L=67\text{ mm}$ driven by an external laser of wavelength $\lambda=1064\text{ nm}$. This configuration ensures a low effective mass of the harmonic oscillator, $m\approx 50\text{ pg}$. The fundamental frequency of the cavity changes with the position $x$ of the membrane as
$\omega(x)=\frac{c}{L}\arccos\bigg{[}|r|\cos\frac{4\pi x}{\lambda}\bigg{]},$ (48)

### v2

@@ IV Results
In this section, we address two main objectives. First, we present results on the system dynamics based on the general solution given in Eq. (32). As previously discussed, we assume the mechanical oscillator is initiated in a thermal state. For the optical field, the initial state is taken to be either
  • (i) a coherent state,
  • (ii) quantum states within the subspace spanned by the vacuum and one-photon Fock states.
Due to the computational challenges associated with evaluating the symmetric logarithmic derivative (SLD), our discussion of the quantum Fisher information (QFI) and classical Fisher information (CFI) will be restricted to the $2\times 2$ photonic subspace, i.e, case (ii).
The results are organized into two main subsections. In subsection IV.1, we explore the dynamics of a dimensionless model in which the coupling parameters $g_{1}$ and $g_{2}$ are treated as independent variables. In subsection IV.2, we consider an experimentally relevant optomechanical setup, where $g_{1}$ and $g_{2}$ are interdependent due to a phenomenological model for $\omega(x)$.
IV.1 Independent couplings
Assuming that $g_{1}$ and $g_{2}$ are independent and can be varied individually, we investigate their respective and combined influences on the system dynamics. To facilitate a clear characterization of the system through dimensionless parameters, we adopt a unitless model defined by
$\hbar=m=\Omega=n_{\text{th}}=\omega_{c}=1.$ (47)
Within this framework, we reveal an intrinsic periodicity in both the system dynamics and the associated QFI/CFI characteristics.
Figure 2: Trajectories in the X–Y plane of the Bloch sphere showing the evolution of the quantum states given in Eqs. (57) and (58), parameterized by $s=0.1$ (red) and $s=0.2$ (blue). The system evolves from $t=0$ to $t=2T_{\omega}$, where $T_{\omega}$ is defined in Eq. (49). The first cycle is shown with solid lines, and the second with dotted lines. The evolution depends solely on the parameter $s$ and is independent of the state’s purity. All other parameters are defined in Eq. 47.
IV.1.1 System dynamics for case (i)
Under realistic conditions, the optical field is often initialized in a coherent states. Thus, we analyze the system dynamics using the Wigner phase-space distribution. This distribution provides a convenient and intuitive way to visualize the state of a quantized radiation field. It is closely related to the Glauber-Sudarshan $P$-representation [21], with the key difference lying in operator ordering: the $P$-representation corresponds to normal ordering of the operators $\hat{a}$ and $\hat{a}^{\dagger}$, while the Wigner phase-space distribution is associated with symmetric ordering. Given the density matrix of a field state $\hat{\rho}$, the Wigner phase-space distribution is given as follows:
$W(\alpha,\alpha^{*})=\frac{1}{\pi^{2}}\int{\rm Tr}\left\{\hat{\rho}\,e^{\zeta\hat{a}^{\dagger}-\zeta^{*}\hat{a}}\right\}e^{\alpha\zeta^{*}-\alpha^{*}\zeta}d^{2}\zeta,$ (48)
where $\alpha,\zeta\in{\mathbb{C}}$.
In Fig. 1, the Wigner phase-space distribution of the optical field is shown, with the field initially prepared in a coherent state of amplitude $\alpha=10$. The Fock space is truncated at $|\alpha|^{2}+4\sqrt{|\alpha|}$, corresponding to a maximum of 140 photons included in the calculations. The figure displays snapshots of the system’s Wigner phase-space distribution at time
$T_{\omega}=\frac{2\pi}{\omega_{1}-\omega_{0}}\quad\text{and}\quad T_{\Omega}=\frac{2\pi}{\Omega_{1}-\Omega_{0}}$ (49)
where $\omega_{n}$ is defined in Eq. 12 and $\Omega_{n}$ in Eq. (7). These times correspond to intermediate characteristic timescales of the system’s evolution, considering the range of frequency differences among all values of $\omega_{n}$ and $\Omega_{n}$. In the figure, the coupling parameter $g_{1}$ increases from left to right, taking the values 0, 11, and $10$, while $g_{2}$ increases from top to bottom. In the Wigner phase-space representation, a coherent state appears as a localized Gaussian packet that undergoes circular motion around the origin in the absence of interaction between the optical field and the mechanical system. However, the interaction between the two oscillators causes the initially localized Gaussian packet to spread along the circular trajectory defined by the free evolution of the optical field. This occurs because the interaction Hamiltonian in Eq. (4) commutes with $\hbar\omega_{c}\hat{a}^{\dagger}\hat{a}$, i.e., with the free evolution of the optical mode. Within this circle, one can observe dense interference fringes. This behavior arises because a coherent state is a specific superposition of Fock states, whose probability amplitudes evolve according to Eq. (32), where $a_{n,m}=\alpha^{n}\alpha^{*m}/\sqrt{n!m!}e^{-|\alpha|^{2}}$. Thus, the time-evolved state of the optical field is a general density matrix of size $140\times 140$. Every density matrix is a convex combination of pure states, which are either Fock states or their superpositions. The convex combination of such states implies only a convex combination of their respective Wigner phase-space distributions. However, due to the large photon number involved up to $n=140$, pure states, unlike coherent states, induce a complex pattern of the interference fringes [21]. The most interesting situation occurs when $g_{2}=0$, because then $T_{\omega}=2\pi/\omega_{c}$. In this situation, the first column of Fig. 1 shows the Wigner function of a coherent state, since the solution of Eq. (36) reduces to:
$a_{n,m}(t)=\frac{\alpha^{n}\alpha^{*m}}{\sqrt{n!m!}}e^{-|\alpha|^{2}}e^{-g_{1}^{2}f^{(2)}_{n,m}(t)-f^{(0)}_{n,m}(t)},$ (50)
where
$f^{(0)}_{n,m}(t)$ $=i\omega_{\text{c}}t(n-m),$ (51)
$f^{(2)}_{n,m}(t)$ $=(2n_{\text{th}}+1)\frac{1-\cos(\Omega t)}{\Omega^{2}}(n-m)^{2}$ (52)
$\qquad-i\frac{\Omega t-\sin(\Omega t)}{\Omega^{2}}(n^{2}-m^{2}).$
Since all parameters in this section are set according to Eq. (47), we have
$f^{(0)}_{n,m}(T_{\omega})$ $=2\pi i(n-m),$ (53)
$f^{(2)}_{n,m}(T_{\omega})$ $=-2\pi i(n^{2}-m^{2}).$ (54)
Furthermore, since $n-m\in\mathbb{Z}$ and $g_{1}$ is chosen as 0, 11, and $10$, it follows that $g^{2}_{1}(n^{2}-m^{2})\in\mathbb{Z}$. Hence,
$a_{n,m}(T_{\omega})$ $=$ $\frac{\alpha^{n}\alpha^{*m}}{\sqrt{n!m!}}e^{-|\alpha|^{2}}e^{-g_{1}^{2}f^{(2)}_{n,m}(T_{\omega})-f^{(0)}_{n,m}(T_{\omega})},$ (55)
$=$ $\frac{\alpha^{n}\alpha^{*m}}{\sqrt{n!m!}}e^{-|\alpha|^{2}}.$
This explains why only a coherent state is observed in those plots.
IV.1.2 System dynamics for case (ii)
@@
To gain deeper insight into the system’s behavior, we consider case (ii), which permits an intuitive visualization of the dynamics on the Bloch sphere. Therefore, in this scenario, the initial photonic state $\hat{\rho}(0)$ is defined on the subspace spanned by the zero-photon state $|0\rangle$ and the one-photon state $|1\rangle$. Thus, $\hat{\rho}(0)$ is a $2\times 2$ positive semi-definite density matrix that can be written as
$\hat{\rho}(0)=\frac{1}{2}(\hat{I}+r_{x}\hat{\sigma}_{x}+r_{y}\hat{\sigma}_{y}+r_{z}\hat{\sigma}_{z}),$ (56)
@@
where $\hat{\sigma}_{i}$ with $i=x,y,z$ are the Pauli matrices and the coefficients $r_{i}$ are real numbers s.t. $r_{x}^{2}+r_{y}^{2}+r_{z}^{2}\leq 1$, with the equality satisfied when $\hat{\rho}(0)$ is a pure state.
In the following discussion, we consider either the pure state
@@
both parametrized by the single parameter $s$. Note that the coefficients $r_{x}$ and $r_{y}$ of the pure states $|\psi_{s}\rangle$ and $|\psi_{1-s}\rangle$ coincide.
In Fig. 2 we show the parametric evolution of the optical field state (36) expressed in terms of the coefficients $r_{x}$ and $r_{y}$, allowing the system evolve between $t=0$ and $t=2T_{\omega}$, defined in Eq. (49). The evolution is shown for two choices of the parameter $s$ of Eqs. (57) and (58). In fact, by solving the integral in Eq. (36), we can see that the diagonal elements of the optical field’s density matrix are time-independent and therefore $r_{z}$ is a constant. Thus, the projection on the X-Y plane of the Bloch sphere suffices to describe the evolution of the optical system, while its value is independent from the purity $\mathcal{P}$ of the system, and can be parametrized solely by $s$. The complex behavior observed in the plots arises for the following reasons. After solving the unitary evolution of the entire system, we traced out the mechanical degrees of freedom without making any approximations. Consequently, the dynamics of the optical field alone reflects the situation in which the two oscillators become entangled and disentangled during the interaction. These processes, in turn, lead to a corresponding decrease or increase in the optical field’s purity. One observes the inward spiral of the coordinates towards the center of the Bloch sphere, which represents the maximally mixed state. Thus, the inward spiral shows a decrease in purity. This happens to a greater degree within the reference time $T_{\omega}$ for larger values of the coupling parameters, as expected from the increased interaction between the optical field and mechanical oscillator.
@@
Figure 3: Time evolution of $r_{x}$, the $x$-component of the system on the Bloch sphere, plotted for the initial pure state $|\psi_{s}\rangle$ with $s=0.2$. Time is expressed in units of the fast oscillation period, two full cycles of which are shown in Fig. 2. The fast oscillation is modulated by a slower beating pattern at frequency $\Delta=\Omega_{1}-\Omega_{0}$. In all subplots, $g_{1}=1.0$ and $g_{2}=0.01$, while the remaining parameters are defined in Eq. 47.
IV.1.3 Spectral Analysis
To demonstrate the system’s approximate periodic behavior, we examine the case where $g_{2}$ is small compared to $g_{1}$, i.e., $g_{1}=1$ and $g_{2}=0.01$. Note that this scenario, where $g_{1}$ is larger than $g_{2}$, is common in experimental setups, as we will present in Sec. IV.2. For this purpose, we show in Fig. 3 the evolution of the $r_{x}$ component. After about $50$ fast cycles, the system spirals into the innermost point corresponding to the lowest purity, and then spirals outward again until it is back at the initial, outermost coordinates corresponding to the highest purity. This process happens at approximately the slower oscillation reference period $T_{\Omega}$, defined in Eq. (49). One can identify the relevant periods from the spectral analysis of $r_{x}$ and purity. We plot the time evolution of the purity $\mathcal{P}$ in Fig. 3 for a pure and mixed state with $s=0.2$.
In order to reveal the underlying frequencies, we performed a Fourier Transform on the dynamics of $r_{x}$ and $\mathcal{P}$, i.e.,
$\tilde{r}_{x}(f)$ $=$ $\int_{\mathbb{R}}dt\ e^{-ift}r_{x}(t),$ (59)
$\tilde{\mathcal{P}}$ $=$ $\int_{\mathbb{R}}dt\ e^{-ift}\mathcal{P}(t).$ (60)
Upon a closer examination of Fig. 3, we find that the primary frequencies of the $r_{x}$ oscillations occur at integer multiples of $\delta/2$, where $\delta=\omega_{1}-\omega_{0}$. This behavior is also evident in Fig. 1, where, in the subplot corresponding to $g_{1}=1$ and $g_{2}=0$, the system completes exactly half a full cycle at $t=T_{\omega}=2\pi/\delta$. While the oscillations and frequencies of $r_{x}$ take place on a shorter timescale, the most prominent periodicity of the system dynamics emerges on a longer timescale, $T_{\Omega}$, defined in Eq. (49) and illustrated in the right-hand subplots of Fig. 3. For both pure and mixed states, the system’s purity oscillates with an almost exact period of $T_{\Omega}$, as further confirmed by the Fourier transform of the purity, which exhibits clear peaks at integer multiples of $\Delta$, where $\Delta=\Omega_{1}-\Omega_{0}$.
Mathematically, the explanation of such periodic behavior of the system can be found in the Gaussian integral that is used to calculate the time dynamics, whose exponent has both real and imaginary components. As the real components may cause damping of the system, the imaginary components cause oscillatory behavior instead. Among all the terms in the integrand exponent, the two frequencies with larger amplitude are $\Delta=\Omega_{1}-\Omega_{0}$, see Eq. (7), and $\delta=\omega_{1}-\omega_{0}$, see Eq. (12). Successive frequencies have an exponentially decaying amplitude and only appear in the spectral analysis as secondary peaks, as shown in Fig. 3.
IV.1.4 Quantum and Classical Fisher Information
In this subsection, we first present, using the setup from case (ii), a comparison between the CFI and QFI for various combinations of $g_{1}$ and $g_{2}$. Our goal is to demonstrate the saturation of the QFI by the CFI. We then examine the QFI properties for case (i). A more detailed analysis is provided in Sec. IV.2, where the parameters of the model are assigned experimentally relevant values.
@@ IV.1.4 Quantum and Classical Fisher Information
Figure 4: Quantum Fisher information as a function of time for the mixed state with $s=0.2$, computed analytically. Periodic peaks appear at regular intervals of $2\pi/\Delta$, with $\Delta=\Omega_{1}-\Omega_{0}$. The parameters are set as $g_{1}=1$, $g_{2}=0.01$, with all other relevant parameters set to unity, as in Eq. 47.
In Fig. 4, we show the temporal evolution of the QFI for the initial mixed state (58) with $s=0.2$. We calculated the QFI using the simplified analytic expression for $\hat{L}_{g}$ given in Equation (43). Even for an initial mixed state, the periodicity of the peaks of the QFI coincide with the periodicity of the purity of the system. Around the peaks, the off-diagonal elements of the density matrix show greater sensitivity to changes of the $g_{2}$ parameter, as can be verified by solving Eq. (36). We also observe that under the same parameter setup, the temporal evolution of the QFI for an initially pure state display much narrower and taller peaks, centered at the same time points as the mixes state counterpart, though this result is not shown here.
Figure 5: Comparison of quantum Fisher information (blue) and classical Fisher information (red) as a function of time for an initially mixed two-level optical field state given by Eq. 58 with $s=0.1$ and for $\Delta=\Omega_{1}-\Omega_{0}$. The time interval spans from 0 to $4T_{\Omega}$, where $T_{\Omega}$ is defined in Eq. (49). All other system parameters are set to unity, as in Eq. 47.
In Fig. 5, various representative combinations of values for $g_{1}$ and $g_{2}$ are taken and the corresponding QFI and CFI evolution in time juxtaposed. In line with the method of BHD outlined in Sec. III, we choose the two-level system in case (ii) to calculate the CFI values for the system. The system for all subplots is initiated in a mixed state labelled by $s=0.1$, see expression 58. As time evolves from 0 to $4T_{\Omega}$, one observes peaks at every regular interval of $T_{\Omega}$. The peaks also increase in intensity as time progresses and decreases in intensity as the values of the coupling parameters increase. Another feature of the correlation is that the peaks of QFI and CFI coincide.
@@
In Fig. 6 we investigate the overlapping peaks of QFI and CFI closely to demonstrate the saturation behavior. By solving Eq. (36), we see that the diagonal elements remain time-independent throughout the system evolution. The set of PVM related to the BHD, as defined in Eq. (46), is chosen to calculate CFI, as it provides information about the off-diagonal elements. When the phase $\phi=0$ they reduce to
$\Pi_{0}=|+\rangle\langle+|,\;\Pi_{1}=|-\rangle\langle-|,$ (61)
@@
where $|\pm\rangle=\frac{1}{\sqrt{2}}(|0\rangle\pm|1\rangle)$, which lead to the measurement probabilities of detecting one or no photon:
$P_{0}=\frac{1}{2}(1+\rho_{01}+\rho_{10}),\;P_{1}=\frac{1}{2}(1-\rho_{01}-\rho_{10}).$ (62)
@@
Note that $\frac{1}{2}(\rho_{01}+\rho_{10})$ is simply the real part of both the off-diagonal elements. We compare the CFI obtained with the above described set of PVM to the QFI in Fig. 6, where we show the result around the slow oscillation reference time $t\approx T_{\Omega}$ in the case of an initial mixed state with $s=0.2$. The CFI has its peak is slightly shifted with respect to the QFI peak. However, we find that when the CFI is maximum, its value corresponds to the QFI value. Similar behavior is obtained both for different choices of $s$ and for an initial pure state.
@@
Figure 6: Classical Fisher information (black dashed line) and quantum Fisher information (red solid line) as a function of time, zoomed in around the first period of the system’s beating oscillation, $t=T_{\Omega}$, where $T_{\Omega}$ is defined in Eq. (49). In this plot, $g_{1}=1.0$, $g_{2}=0.01$, $\phi=\pi/2$, and all other system parameters are set to unity, as in Eq. 47.
Finally, we turn to case (i) and demonstrate the periodic behavior of QFI for an optical system initialized in a coherent state with $\alpha=3$. As the system size has increased, the analytical two-level system solution to QFI no longer applies, hence numerical integration method is adopted. QFI as a function of time is presented as a scatter plot, see Fig. 7. The transparency of the data series’ coloring is taken as a linear function calculated as the following:
$\alpha=\frac{Q\times T^{2}_{\Omega}}{t^{2}\times\max(\frac{Q\times T^{2}_{\Omega}}{t^{2}})},$ (63)
such that the transparency parameter $\alpha$ is a distribution on the interval $[0,1].$
Figure 7: Quantum Fisher information against time as a scatter plot with $\Delta=\Omega_{1}-\Omega_{0}$. One observes peaks at regular interval with increasing magnitude. The system is initialized in a coherent state with $\alpha=3$, with all system parameters set to unity as in Eq. 47, while $g_{1}=1.0$ and $g_{2}=0.01$. The transparency of the data series is scaled by expression (63). A quadratic relation over the peaks is plotted for reference with $Q\approx 4.334\times 10^{11}\times\frac{t^{2}}{T_{\Omega}^{2}}$.
Furthermore, a quadratic data series is juxtaposed on the scatter plot to fit the successive peaks observed, it follows the equation:
$y=4.334\times 10^{11}\times\frac{t^{2}}{T_{\Omega}^{2}}.$
From this, one sees that the intensity of the peaks scales quadratically with system time. This can be explained by the increasing amount of interaction between the optical field and the mechanical oscillator. Given that up to quadratic interaction is considered, a quadratic increase in the extractable information can be expected.
IV.2 Interdependent couplings
In the preceding sections, $g_{1}$ and $g_{2}$ were treated as independent variables. In a laboratory setting, however, certain phenomenological models approximate the cavity frequency as a function of some of the system parameters. In the following subsection, we show that such a model implies a mathematical relation between $g_{1}$ and $g_{2}$ when Eq. (2) is truncated strictly at the quadratic term; that is, all terms of order higher than two are assumed to have zero coefficients in the Taylor expansion. For the results presented below, we adopt this strict approximation and derive the corresponding relation between $g_{1}$ and $g_{2}$.
We consider the following phenomenological model [37]:
$\omega(x)=\frac{c}{L}\arccos\bigg[r\cos\frac{4\pi x}{\lambda}\bigg],$ (64)
where $L$ denotes the cavity length, $r$ the membrane reflectivity, $\lambda$ the wavelength of the laser driving the cavity, and $x$ is the membrane center-of-mass position along the cavity axis [7, 14]. By combining Eq. (2) with Eq. (64), an interdependent relation between $g_{1}$ and $g_{2}$ can be established. More specifically,
$\omega^{\prime}(x)=\frac{cr}{L}\frac{\frac{4\pi}{\lambda}\sin\left(\frac{4\pi x}{\lambda}\right)}{\sqrt{1-\left[r\cos\left(\frac{4\pi x}{\lambda}\right)\right]^{2}}},$ (65)
which has zeros at $x=(n\lambda)/2$, where $n$ is an integer. From Eq. 2, we obtain
$=$ $\left[\omega(x_{0})+\frac{\omega^{\prime\prime}(x_{0})}{2!}x_{0}^{2}\right]+\left[\omega^{\prime}(x_{0})-\omega^{\prime\prime}(x_{0})x_{0}\right]x$
$+$ $\frac{\omega^{\prime\prime}(x_{0})}{2!}x^{2}=\omega_{c}+g_{1}x+g_{2}x^{2},$
To have $\omega^{\prime}(x_{0})=0$, we set $x_{0}=(n\lambda)/2$ with $n>0$, which yields the following relations
$\omega_{c}$ $=$ $\omega(x_{0})+\frac{\omega^{\prime\prime}(x_{0})}{2!}x_{0}^{2},$ (67)
$g_{1}$ $=$ $-\omega^{\prime\prime}(x_{0})x_{0},$ (68)
$g_{2}$ $=$ $\frac{\omega^{\prime\prime}(x_{0})}{2!}.$ (69)
Reference Values
$\hbar$ $1.054\times 10^{-34}$ $\text{J}\cdot\text{s}]$
$m$ $50\times 10^{-15}$ $\text{kg}]$
$\Omega$ $2\pi\times 134\times 10^{3}$ $\text{Hz}]$
$n_{th}$ $4.665\times 10^{7}$
$\omega_{c}$ $6.544\times 10^{15}$ $\text{Hz}]$
$g_{1}$ $1.308\times 10^{18}$ $\text{Hz}/\text{m}]$
$g_{2}$ $1.445\times 10^{23}$ $\text{Hz}/\text{m}^{2}]$
Table 1: Table of parameter values used as a reference for an experimental system. All parameter values are expressed in SI units. Among the listed parameters, $m$ and $\Omega$ are directly taken from [37], while $n_{\text{th}}$, the average phonon number, is calculated for a temperature of $300,\text{K}$. The remaining parameters are obtained by substituting the experimental values of Ref. [37] required for the phenomenological model in Eq. (64) into Eqs. (67), (68), and (69).